OpenStudy (anonymous):

-p^2+30p How do I graph that? I don't know the y-int

OpenStudy (phi):

You know it's a parabola, so it's U shaped or upside down U. -p^2 means upside down. factor out p: y= p * -(p+30) = p*(30-p) y is 0 when p=0 or p= 30 it peaks halfway between 0 and 30, or at p= 15. plug p=15 into the equation to find y

OpenStudy (turingtest):

find the y-int by setting x=0. you get (0,0) or y-int=0. like phi said it is an inverted parabola stretching from (0,0) to (30,0). I am assuming p is on the x-axis, right?

OpenStudy (turingtest):

by the way the vertex is (15,225)